OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = −2

OpenStudy (anonymous):

y=1-2x 4x+2(1-2x)=-2 -2x=-4 x=2 2(2) + y = 1 y+-3

OpenStudy (anonymous):

Solve the first equation for Y. Then substitute that value into the second equation. If you show your work here we can help to show you what's right and what's wrong.

OpenStudy (anonymous):

a. x = 0 b. x = 2 c. There is no x value as there is no solution. d. x can be any value as there is an infinite number of solutions.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

B. x=2

OpenStudy (anonymous):

Thanks <3

OpenStudy (anonymous):

Welcome! And my first post is supposed to say y=-3 at the end not y+-3